# Designing Concise ConvNets with Columnar Stages

### v1

@@ 2 Related Work
This section provides an overview of representative network designs (Figure 1). The earlier ConvNets Krizhevsky et al. (2012); Simonyan & Zisserman (2014) stacked dense convolutions with an increasing number of channels and decreasing resolution (Figure 1a). Improved versions He et al. (2016); Szegedy et al. (2015); Xie et al. (2017) achieve higher accuracy via manually designed blocks (Figure 1c), while Howard et al. (2017); Ma et al. (2018); Sandler et al. (2018); Zhang et al. (2018), use depthwise convolutions Sifre & Mallat for saving computations.
ConvNets have also grown from branchless Krizhevsky et al. (2012); Simonyan & Zisserman (2014) (Figure 1a) to single branch He et al. (2016) (Figure 1c) to multi-branch Radosavovic et al. (2020); Szegedy et al. (2016); Tan & Le (2019); Zoph et al. (2018) (Figure 1b). These models utilize $1\times 1$ convolutions frequently, which rapidly increases network depth He et al. (2016); Sandler et al. (2018); Tan & Le (2019); Zhang et al. (2018) (Figure 1c-1e). Although beneficial, both large depth and high branching tend to increase the latency, memory requirements, and Memory Access Cost (MAC) Chen et al. (2024) due to the serialized execution of parallel branches Ding et al. (2021); Srivastava et al. (2015); Tan & Le (2019).
Recent RepVGG Ding et al. (2021) proposes structural parameterization (SR) to resolve the branching issue. While ParNet Goyal et al. (2021) and VanillaNet Chen et al. (2024) reduce depth to achieve lower latency. Efforts to reduce depth increase the parameter count Chen et al. (2024); Goyal et al. (2021) to match the accuracy of relatively deeper counterparts He et al. (2016).
Recent Vision Transformers (ViTs) Dosovitskiy et al. (2020); Liu et al. (2023; 2021); Touvron et al. (2021) have attracted huge research interests. As outlined in Dai et al. (2021b), the $O(N^{2})$-complex attention in ViTs is a notable issue from a data size and resource-constrained viewpoint. This issue continues to inspire improvements in ConvNets. For instance, RepLKNet Ding et al. (2022) aims to bridge the gap between ViT and CNNs by employing large kernels.
The previously discussed designs focus on particular aspects, e.g., He et al. (2016); Xie et al. (2017) on the accuracy, Chen et al. (2024); Goyal et al. (2021) on runtime and depth. On the other hand, our aim is not to develop large-scale complex models. Instead, given the success of VanillaNet-style networks, we focus on our sub-objectives (Sec 1) to push the frontier of simple, concise models by revisiting the representative ConvNets.
@@ 3 Columnar Stage Network
Figure 2: Design evolution flow of CoSNet-unit. (a) A ResNet He et al. (2016) stage with three blocks. (b) removing all $1\times 1$ convolutions except the first of the first block and the last of the last block. (c) detailed design of the CoSNet-unit by integrating our design ideas into ‘(b)’, and (d) final optimized CoSNet-unit from an implementation viewpoint.
Our approach is a series of improvements motivated by representative ConvNet designs. To understand better, we begin with ResNet He et al. (2016) as a stepping stone as done in Liu et al. (2022). We design the building block of CoSNet i.e., CoSNet-unit while recalling our sub-objectives: 1)1)1 ) reducing depth, 2)2)2 ) controlled parameter count, 3)3)3 ) high computational density, 4)4)4 ) uniform primitive operation, and 5)5)5 ) low branching.
@@ 3.1 Avoiding $1\times 1$ for Reducing Depth
The recent works of reducing depth Chen et al. (2024); Goyal et al. (2021) increase the parameter count to achieve accuracy similar to a deeper network. However, we aim to reduce depth while avoiding a large parameter count, which is a difficult objective. Hence, we handle reducing depth and controlling parameter count separately.
To reduce depth, we identify that $1\times 1$ convolutions in the ResNet-like designs (Figure 2a) (He et al., 2016; Liu et al., 2022) etc., form almost $66\%$ of depth without improving receptive field due to their pointwise nature Luo et al. (2016). Hence, we minimize the number of these layers. Specifically, we use only two $1\times 1$ convolutions $L_{s}$ and $L_{f}$ in a CoSNet-unit, where $L_{s}$ reduces the channel squeezing while $L_{f}$ performs expansion (Figure 2b). Then, we stack $l$ number of $3\times 3$ convolutions, forming a column sandwiched between $L_{s}$ and $L_{f}$.
@@
We propose Parallel Columnar Convolutions to handle the large parameter count originating to compensate for the lost non-linearity due to the reduced depth Chen et al. (2024). In this design, we first deploy $M$ columns in parallel (Figure 2c), and crosstalk among columns does not exist, i.e. a convolution of a column can only feed a convolution of the same column. Then, we restrict the number of kernels in a convolution layer of a column to a small number of $N$. This design affects the number of parameters less aggressively when the number of columns increases (see ablations in the supplement). This is a powerful feature of CoSNet design, offering controlled growth of parameter count during network scaling. This helps CoSNet achieve higher accuracy with fewer parameters.
The idea of the parallel column is based on our hypothesis that multiple kernels with fewer channels can be better than one with large channels. Having $M$ convolutions in parallel with a smaller number of kernels $N$ is equivalent to synthesizing multiple kernels from a large kernel. On the other hand, the idea of smaller $N$ is motivated by the fact that many parallelly operating neurons tend to learn redundant representations while being computationally taxing and causing overfitting. For the same reason, EfficientViT Liu et al. (2023) slices the input channels in its structure. Hence, by keeping $N$ small, we expect to decouple the data patterns learned by the different columns.
In ConvNets, a similar idea was proposed in Inception Szegedy et al. (2015), then in ResNeXt Xie et al. (2017), and then abandoned later as it caused inefficiency. For instance, Inception uses different-sized convolutions and pooling in parallel, which must be executed serially despite being employed in parallel. Also, Inception differs from our columnar architecture since it does not have columns as deep as CoSNet.
@@ 3.3 Input Replication
Input replication has also been employed in the earlier ResNeXt Xie et al. (2017), but notable differences exist. ResNeXt has multiple blocks per stage, and each block performs IR, as shown in Figure 1d. Whereas CoSNet performs IR only once. In ResNeXt, IR is performed before $1\times 1$ squeeze layer, whereas in CoSNet, it is done after the squeeze layer.
The parallel columnar organization may seem to overlap with widely explored group convolutions Xie et al. (2017); Zhang et al. (2018). However, there are two key differences. First, group convolution divides the input channels, thus defying the objective of IR because now each column receives only a subset of the input channels, thus less information per group, as shown in Figure 1g. On the contrary, CoSNet uses IR, which feeds each column with the replica of the input, thus making the entire input information accessible to each column. This becomes one of the reasons that despite infrequent fusion (Sec. 3.6), unlike group conv, CoSNet still performs better (See ablations in the supplement).
@@
With this optimization, all columns can be efficiently processed using GPU-based highly optimized Batched-Matrix-Multiply routines, leading to increased computational density, increased GPU utilization, reduced memory access cost Ding et al. (2021), and minimized GPU load-dispatch transactions. Thus resulting in a simplified CoSNet design (Figure 2d). Moreover, since an CoSNet-unit is made up mostly of $3\times 3$ convolutions, it well suits the convolution hardware accelerators because they have dedicated support for them, and more chip area can be dedicated to $3\times 3$ computational units.
@@
From the previous step, batched processing yields additional benefits, i.e., CoSNet becomes uni-branched regardless of training and testing. This reduces memory consumption and access costs, resulting in lower per-iteration training time and increased parallelization. This contrasts with RepVGG Ding et al. (2021), which has a considerable training time. Regarding ASIC development, low branching in CoSNet leaves more area on the chip because of the reduced memory requirement to store intermediate tensors. This area can now be dedicated to more computational units.
Although the multi-branch design is beneficial for achieving high accuracy Ding et al. (2021) (Figure 1f), CoSNet, despite having minimal branching, effortlessly achieves high accuracy. This is because the core design of CoSNet-unit posses multiple branches in the form of columns and short projections (Figure 2c). However, due to batched processing CoSNet-unit mimics uni-branched behavior. In this way, CoSNet takes advantage of both worlds, i.e., eliminated train time complexity due to multiple branches and fast inference during test time without needing structural parameterization Ding et al. (2021).
@@ 3.6 Fuse Once
Finally, the output of all the columns is fused by $L_{f}$. In ResNeXt (Figure 1d), the output of $3\times 3$ convs are fused immediately via a $1\times 1$ conv, whereas in CoSNet, it is done much later. Our fuse once strategy is different from group Zhang et al. (2018) or depthwise convolutions Howard et al. (2017) that are followed by $1\times 1$ (Figure 1g) to avoid loss of accuracy because each group/channel has too few connections which restrict its learning ability without frequent fusion (Zhang et al. (2018), Figure 1g). This increases network depth and, hence, latency. On the contrary, CoSNet is free from this constraint because we increase $N$ as we go deep in CoSNet unlike Zhang et al. (2018). Hence, each neuron in $M$ columns has a sufficiently large number of connections that enable learning without frequent fusion. We performed an ablation (see supplement) by applying the same strategy as Figure 1g in CoSNet. We observed increased network depth, latency, and decreased accuracy.
Figure 3: Illustration of Vanilla Frequent Fusion (left) (Zhang et al. (2018), Figure 1g) and Pairwise Frequent Fusion (right).
@@ 3.7 Projections
To facilitate better gradient flow during network training, we employ projections introduced by ResNet He et al. (2016) but slightly differently in two ways:
@@
The above projection design helps achieve CoSNet better accuracy (see ablations) and is slightly different from the existing ones. First, projection in ResNet-like models He et al. (2016); Xie et al. (2017) is used only in the first block of a stage (shallower), and projection between stages does not exist. Second, projection in these models operates at a stride of $2$. On the contrary, in CoSNet, the projection connects two stages (deeper) while operating at unit stride and utilizing pooling to increase the receptive field.
@@ 3.8 CoSNet Instantiation
A CoSNet variant can be instantiated by stacking CoSNet-units (Figure 4). CoSNet does not have the notion of blocks but only has stages in the form of CoSNet-unit. This contrasts with existing ConvNets, which have stages, and each stage comprises multiple blocks Goyal et al. (2021); He et al. (2016); Liu et al. (2022); Xie et al. (2017) e.g., ResNet-$50$ has four stages, having $3$, $4$, $6$, and $3$ blocks respectively (Figure 1c1e).
To instantiate a CoSNet variant, we follow the tradition of five stages He et al. (2016); Simonyan & Zisserman (2014), among which the first (stem) is a $3\times 3$ convolution with a stride of $2$, while the remaining are the CoSNet-units.
Following ResNet He et al. (2016), we set channels of $L_{s}$ to $64$, which gets doubled at each stage, while the channels of $L_{p}$ and $L_{f}$ always equal to $\zeta$ times the channels of $L_{s}$. We set $\zeta=4$, following He et al. (2016). To further simplify the instantiation, we set the depth of a column, i.e. $l$ in $k^{th}$ CoSNet-unit equal to the number of blocks in the $k^{th}$ stage of a widely used model ResNet-$50$ He et al. (2016). Summarily, CoSNet-unit has only three hyperparameters: $M,N,l$ which can be changed to construct different CoSNet variant. Please refer to the supplement for CoSNet instance names.
@@ 4 Experiments
We evaluate CoSNet on ImageNet Deng et al. (2009) dataset consisting of $1.28$M train and $50$k validation images of $1000$ categories. Our training methodology is consistent with recent VanillaNet Chen et al. (2024). We use data augmentation techniques in Chen et al. (2024); Liu et al. (2022). See the appendix at the end of this paper for more details.
@@ 4.1 Advanced ConvNets and Vision Transformers
CoSNet vs recent EfficientViT Liu et al. (2023) As shown in Table 1 and Figure 5, CoSNet is less deep and runs $60\%$ faster than EfficientViT Transformer while exhibiting better accuracy, e.g., EfficientVit-M4444 vs CoSNet-A00. EfficientVit is another example of lower FLOPs that do not guarantee lower latency. Even the CoSNet-A1111-PFF variant is still relatively shallower than EfficientVit while delivering better accuracy.
CoSNet vs DeiT Touvron et al. (2021) From Table 1, CoSNet-B1111 is almost $50\%$ less deep, has $23\%$ fewer params, and runs $60\%$ faster than DeiT Transformer while exhibiting slightly better accuracy. With PFF, CoSNet-B00-PFF performs better in terms of accuracy, depth, and runtime.
CoSNet vs advanced mid-range ConvNets and Transformers CoSNet-B2222 is $72\%$ less deeper, $55\%$ faster, and $1.2\%$ more accurate than the popular Swin Transformer Liu et al. (2021). It is also $55\%$ less deeper, $30\%$ faster with slightly lower accuracy than the popular ConvNeXt Liu et al. (2022). Moreover, CoSNet-C2222 rivals the latest ConvNext-v2222-T Woo et al. (2023) with similar accuracy but higher speed and smaller depth.
CoSNet-B2222, C1111, C2222 models rivals advanced Transformers, such as ViTAE-S Xu et al. (2021) and hybrid models, such as CoAtNet-0 Dai et al. (2021b). With similar parameter counts and accuracy, our models show faster inference speed. The competitive tradeoffs offered by CoSNet show the significance of concise models.

### v2

@@ 2 Related Work
This section provides an overview of representative network designs (Figure 1). The earlier ConvNets (Krizhevsky et al., 2012; Simonyan & Zisserman, 2014) stacked dense convolutions with an increasing number of channels and decreasing resolution (Figure 1a). Improved versions (He et al., 2016; Szegedy et al., 2015; Xie et al., 2017) achieve higher accuracy via manually designed blocks (Figure 1c), while (Howard et al., 2017; Ma et al., 2018; Sandler et al., 2018; Zhang et al., 2018), use depthwise convolutions (Sifre & Mallat, ) for saving computations, but they are not memory friendly (Ding et al., 2021).
ConvNets have also grown from branchless (Krizhevsky et al., 2012; Simonyan & Zisserman, 2014) (Figure 1a) to single branch (He et al., 2016) (Figure 1c) to multi-branch (Radosavovic et al., 2020; Szegedy et al., 2016; Tan & Le, 2019; Zoph et al., 2018) (Figure 1b). These models utilize $1\times 1$ convolutions frequently, which rapidly increases network depth (He et al., 2016; Sandler et al., 2018; Tan & Le, 2019; Zhang et al., 2018) (Figure 1c-1e). Although beneficial, both large depth and high branching tend to increase the latency, memory requirements, and Memory Access Cost (MAC) (Chen et al., 2023) due to the serialized execution of parallel branches (Ding et al., 2021; Srivastava et al., 2015; Tan & Le, 2019).
Recent RepVGG (Ding et al., 2021) proposes structural parameterization (SR) to resolve the branching issue. While ParNet (Goyal et al., 2021) and VanillaNet (Chen et al., 2023) reduce depth to achieve lower latency. Efforts to reduce depth increase the parameter count (Chen et al., 2023; Goyal et al., 2021) to match the accuracy of relatively deeper counterparts (He et al., 2016).
Recent Vision Transformers (ViTs) (Dosovitskiy et al., 2020; Liu et al., 2023; 2021; Touvron et al., 2021) have attracted huge research interests. As outlined in (Dai et al., 2021b), the $O(N^{2})$-complex attention in ViTs is a notable issue from a data size and resource-constrained viewpoint. This issue continues to inspire improvements in ConvNets. For instance, RepLKNet (Ding et al., 2022) aims to bridge the gap between ViT and CNNs by employing large kernels.
The above designs focus on limited aspects, e.g., (He et al., 2016; Xie et al., 2017) on the accuracy, (Chen et al., 2023; Goyal et al., 2021) on runtime and depth. To address this research gap, we draw inspiration from the success of VanillaNet-style networks, and instead of pursuing large-scale models, we focus on our sub-objectives (Sec 1) and revisit the representative ConvNets to push the frontier of simple, concise models.
@@ 3 Columnar Stage Network
Figure 2: Design evolution flow of CoSNet-unit. (a) A ResNet (He et al., 2016) stage with three blocks. (b) removing all $1\times 1$ convolutions except the first of the first block and the last of the last block. (c) detailed design of the CoSNet-unit by integrating our design ideas into ‘(b)’, and (d) final optimized CoSNet-unit from an implementation viewpoint.
Our approach is a series of improvements motivated by representative ConvNet designs. To understand better, we begin with ResNet (He et al., 2016) as a stepping stone as done in (Liu et al., 2022). We design the building block of CoSNet i.e., CoSNet-unit while recalling our sub-objectives: 1)1)1 ) reducing depth, 2)2)2 ) controlled parameter count, 3)3)3 ) high computational density, 4)4)4 ) uniform primitive operation, and 5)5)5 ) low branching.
@@ 3.1 Avoiding $1\times 1$ for Reducing Depth
The recent works of reducing depth (Chen et al., 2023; Goyal et al., 2021) increase the parameter count to achieve accuracy similar to a deeper network. However, we aim to reduce depth while avoiding a large parameter count, which is a difficult objective. Hence, we handle reducing depth and controlling parameter count separately.
To reduce depth, we identify that $1\times 1$ convolutions in the ResNet-like designs (Figure 2a) (He et al., 2016; Liu et al., 2022) etc., form almost $66\%$ of depth without improving receptive field due to their pointwise nature (Luo et al., 2016). Hence, we minimize the number of these layers. Specifically, we use only two $1\times 1$ convolutions $L_{s}$ and $L_{f}$ in a CoSNet-unit, where $L_{s}$ reduces the channel squeezing while $L_{f}$ performs expansion (Figure 2b). Then, we stack $l$ number of $3\times 3$ convolutions, forming a column sandwiched between $L_{s}$ and $L_{f}$.
@@
We propose Parallel Columnar Convolutions to handle the large parameter count originating to compensate for the lost non-linearity due to the reduced depth (Chen et al., 2023). In this design, we first deploy $M$ columns in parallel (Figure 2c), and crosstalk among columns does not exist, i.e. a convolution of a column can only feed a convolution of the same column. Then, we restrict the number of kernels in a convolution layer of a column to a small number of $N$. This design affects the number of parameters less aggressively when the number of columns increases (see ablations in the supplement). This is a powerful feature of CoSNet design, offering controlled growth of parameter count during network scaling. This helps CoSNet achieve higher accuracy with fewer parameters.
The idea of the parallel column is based on our hypothesis that multiple kernels with fewer channels can be better than one with large channels. Having $M$ convolutions in parallel with a smaller number of kernels $N$ is equivalent to synthesizing multiple kernels from a large kernel. On the other hand, the idea of smaller $N$ is motivated by the fact that many parallelly operating neurons tend to learn redundant representations while being computationally taxing and causing overfitting. For the same reason, EfficientViT (Liu et al., 2023) slices the input channels in its structure. Hence, by keeping $N$ small, we expect to decouple the data patterns learned by the different columns.
In ConvNets, a similar idea was proposed in Inception (Szegedy et al., 2015), then in ResNeXt (Xie et al., 2017), and then abandoned later as it caused inefficiency. For instance, Inception uses different-sized convolutions and pooling in parallel, which must be executed serially despite being employed in parallel. Also, Inception differs from our columnar architecture since it does not have columns as deep as CoSNet.
@@ 3.3 Input Replication
Input replication has also been employed in the earlier ResNeXt (Xie et al., 2017), but notable differences exist. ResNeXt has multiple blocks per stage, and each block performs IR, as shown in Figure 1d. Whereas CoSNet performs IR only once. In ResNeXt, IR is performed before $1\times 1$ squeeze layer, whereas in CoSNet, it is done after the squeeze layer.
The parallel columnar organization may seem to overlap with widely explored group convolutions (Xie et al., 2017; Zhang et al., 2018). However, there are two key differences. First, group convolution divides the input channels, thus defying the objective of IR because now each column receives only a subset of the input channels, thus less information per group, as shown in Figure 1g. On the contrary, CoSNet uses IR, which feeds each column with the replica of the input, thus making the entire input information accessible to each column. This becomes one of the reasons that despite infrequent fusion (Sec. 3.6), unlike group conv, CoSNet still performs better (See ablations in the supplement).
@@
With this optimization, all columns (Figure 2c) can be efficiently processed using GPU-based highly optimized Batched-Matrix-Multiply routines, leading to increased computational density, increased GPU utilization, reduced memory access cost (Ding et al., 2021), and minimized GPU load-dispatch transactions. Thus resulting in a simplified CoSNet design (Figure 2d). Moreover, since an CoSNet-unit is made up mostly of $3\times 3$ convolutions, it well suits the convolution hardware accelerators because they have dedicated support for them, and more chip area can be dedicated to $3\times 3$ computational units.
@@
From the previous step, batched processing yields additional benefits, i.e., CoSNet becomes uni-branched regardless of training and testing. This reduces memory consumption and access costs, resulting in lower per-iteration training time and increased parallelization. This contrasts with RepVGG (Ding et al., 2021), which has a considerable training time. Regarding ASIC development, low branching in CoSNet leaves more area on the chip because of the reduced memory requirement to store intermediate tensors. This area can now be dedicated to more computational units.
Although the multi-branch design is beneficial for achieving high accuracy (Ding et al., 2021) (Figure 1f), CoSNet, despite having minimal branching, effortlessly achieves high accuracy. This is because the core design of CoSNet-unit posses multiple branches in the form of columns and short projections (Figure 2c). However, due to batched processing CoSNet-unit mimics uni-branched behavior. In this way, CoSNet takes advantage of both worlds, i.e., eliminated train time complexity due to multiple branches and fast inference during test time without needing structural parameterization (Ding et al., 2021).
@@ 3.6 Fuse Once
Finally, the output of all the columns is fused by $L_{f}$. In ResNeXt (Figure 1d), the output of $3\times 3$ convs are fused immediately via a $1\times 1$ conv, whereas in CoSNet, it is done much later. Our fuse once strategy is different from group (Zhang et al., 2018) or depthwise convolutions (Howard et al., 2017) that are followed by $1\times 1$ (Figure 1g) to avoid loss of accuracy because each group/channel has too few connections which restrict its learning ability without frequent fusion ((Zhang et al., 2018), Figure 1g). This increases network depth and, hence, latency. On the contrary, CoSNet is free from this constraint because we increase $N$ as we go deep in CoSNet unlike (Zhang et al., 2018). Hence, each neuron in $M$ columns has a sufficiently large number of connections that enable learning without frequent fusion. We performed an ablation (see supplement) by applying the same strategy as Figure 1g in CoSNet. We observed increased network depth, latency, and decreased accuracy.
Figure 3: Illustration of Vanilla Frequent Fusion (left) ((Zhang et al., 2018), Figure 1g) and Pairwise Frequent Fusion (right).
@@ 3.7 Projections
To facilitate better gradient flow during network training, we employ projections introduced by ResNet (He et al., 2016) but slightly differently in two ways:
@@
The above projection design helps achieve CoSNet better accuracy (see ablations) and is slightly different from the existing ones. First, projection in ResNet-like models (He et al., 2016; Xie et al., 2017) is used only in the first block of a stage (shallower), and projection between stages does not exist. Second, projection in these models operates at a stride of $2$. On the contrary, in CoSNet, the projection connects two stages (deeper) while operating at unit stride and utilizing pooling to increase the receptive field.
@@ 3.8 CoSNet Instantiation
A CoSNet variant can be instantiated by stacking CoSNet-units (Figure 4). CoSNet does not have the notion of blocks but only has stages in the form of CoSNet-unit. This contrasts with existing ConvNets, which have stages, and each stage comprises multiple blocks (Goyal et al., 2021; He et al., 2016; Liu et al., 2022; Xie et al., 2017) e.g., ResNet-$50$ has four stages, having $3$, $4$, $6$, and $3$ blocks respectively (Figure 1c1e).
To instantiate a CoSNet variant, we follow the tradition of five stages (He et al., 2016; Simonyan & Zisserman, 2014), among which the first (stem) is a $3\times 3$ convolution with a stride of $2$, while the remaining are the CoSNet-units.
Following ResNet (He et al., 2016), we set channels of $L_{s}$ to $64$, which gets doubled at each stage, while the channels of $L_{p}$ and $L_{f}$ always equal to $\zeta$ times the channels of $L_{s}$. We set $\zeta=4$, following (He et al., 2016). To further simplify the instantiation, we set the depth of a column, i.e., $l$ in $k^{th}$ CoSNet-unit equal to the number of blocks in the $k^{th}$ stage of a widely used model ResNet-$50$ (He et al., 2016). Summarily, CoSNet-unit has only three hyperparameters: $M,N,l$ which control CoSNet’s parameters, depth, latency, and accuracy. Hence, different CoSNet variants can be constructed by changing them. Please refer to the supplement for CoSNet instance names and ablations on $M,N,l$.
@@ 4 Experiments
We evaluate CoSNet on ImageNet (Deng et al., 2009) dataset consisting of $1.28$M train and $50$k validation images of $1000$ categories. Our training methodology is consistent with recent VanillaNet (Chen et al., 2023). We use data augmentation techniques in (Chen et al., 2023; Liu et al., 2022). See the appendix at the end of this paper for more details.
@@
$\bullet$ InceptionNeXt-S (Yu et al., 2024) ConvNet $48$ $49.0$M $8.40$B $18$ms $55$ $83.5$
$\bullet$ UniRepLKNet-S (Ding et al., 2024) ConvNet $180$ $56.0$M $9.10$B $23$ms $43$ $83.9$
@@ 4.1 Advanced ConvNets and Vision Transformers
CoSNet vs recent EfficientViT (Liu et al., 2023) As shown in Table 1 and Figure 5, CoSNet is less deep and runs $60\%$ faster than EfficientViT Transformer while exhibiting better accuracy, e.g., EfficientVit-M4444 vs CoSNet-A00. EfficientVit is another example of lower FLOPs that do not guarantee lower latency. Even the CoSNet-A1111-PFF variant is still relatively shallower than EfficientVit while delivering better accuracy.
CoSNet vs DeiT (Touvron et al., 2021) From Table 1, CoSNet-B1111 is almost $50\%$ less deep, has $23\%$ fewer params, and runs $60\%$ faster than DeiT Transformer while exhibiting slightly better accuracy. With PFF, CoSNet-B00-PFF performs better in terms of accuracy, depth, and runtime.
CoSNet vs advanced mid-range ConvNets and Transformers CoSNet-B2222 is $72\%$ less deeper, $55\%$ faster, and $1.2\%$ more accurate than the popular Swin Transformer (Liu et al., 2021). It is also $55\%$ less deeper, $30\%$ faster with slightly lower accuracy than the popular ConvNeXt (Liu et al., 2022). Moreover, CoSNet-C2222 rivals the latest ConvNext-v2222-T (Woo et al., 2023) with similar accuracy but higher speed and smaller depth.
CoSNet-B2222, C1111, C2222 models rivals advanced Transformers, such as ViTAE-S (Xu et al., 2021) and hybrid models, such as CoAtNet-0 (Dai et al., 2021b). With similar parameter counts and accuracy, our models show faster inference speed. The competitive tradeoffs offered by CoSNet show the significance of concise models.